OpenStudy (ashleynguyenx3):

Could you help me find the derivative of this logarithmic function? (problem below)

OpenStudy (ashleynguyenx3):

\[g(x)=\int\limits_{1}^{\ln x}(t^2+3)dt\]

OpenStudy (anonymous):

ummmm it looks like an integral!

OpenStudy (ashleynguyenx3):

Yeah, I know. But it said to find the derivative.

OpenStudy (ashleynguyenx3):

So would I just plug in lnx and 1?

OpenStudy (anonymous):

shouldn't it be \[f(t)\int\limits_{1}^{lnt}(t ^{2}+3)dt\]

OpenStudy (ashleynguyenx3):

Yeah, sorry.

OpenStudy (ashleynguyenx3):

It's just a variable.

OpenStudy (ashleynguyenx3):

So do I just use the second fundamental theorem of calculus?

OpenStudy (anonymous):

with this notation yes......i don't understand why they would tell you to find the dervative with integral notation

OpenStudy (ashleynguyenx3):

I don't know, but thanks. :D

OpenStudy (anonymous):

no prob